OpenStudy (anonymous):

write the given expression in algebraic form: tan(arccos(x/3))

zepdrix (zepdrix):

\[\Large \tan\left(\color{royalblue}{\arccos\frac{x}{3}}\right)\]So let's deal with the blue part first. We need to make a triangle with it.\[\Large \color{royalblue}{\arccos\frac{x}{3}=\theta} \qquad\to\qquad \cos \theta=\frac{x}{3}\]

zepdrix (zepdrix):

\[\Large \cos \theta=\frac{adjacent}{hypotenuse}\]

zepdrix (zepdrix):



zepdrix (zepdrix):

\[\Large \tan\left(\color{royalblue}{\arccos\frac{x}{3}}\right) \qquad\to\qquad \tan\left(\color{royalblue}{\theta}\right)=\frac{opposite}{adjacent}\]

zepdrix (zepdrix):

Gotta find that missing side